OpenStudy (anonymous):

Solve 7-7x=(3x+2)(x-1) and tell whether or not it is extraneous

OpenStudy (anonymous):

7-7x=3x^2-3x+2x-2 3x^2+6x-9=0 3(x^2+2x-3)=0 3(x-1)(x+3)=0 the answers are 1 and -3.